# Entropy Sentinel: Continuous LLM Accuracy Monitoring from Decoding Entropy Traces in STEM

### v2

@@ Abstract
Deploying LLMs raises two coupled challenges: (1) monitoring—estimating where a model underperforms as traffic and domains drift—and (2) improvement—prioritizing data acquisition to close the largest performance gaps. We test whether an inference-time signal can estimate slice-level accuracy under domain shift. For each response, we compute an output-entropy profile from final-layer next-token probabilities (from top-$k$ logprobs) and summarize it with eleven statistics. A lightweight classifier predicts instance correctness, and averaging predicted probabilities yields a domain-level accuracy estimate. We evaluate on ten STEM reasoning benchmarks with exhaustive train/test compositions ($k\in\{1,2,3,4\}$; all $\binom{10}{k}$ combinations), across nine LLMs from six families (3B–20B). Estimates often track held-out benchmark accuracy, and several models show near-monotonic ordering of domains. Output-entropy profiles are thus an accessible signal for scalable monitoring and for targeting data acquisition.
@@
Pedro Memoli Buffa Departamento de Matematica, FCEyN Universidad de Buenos Aires Buenos Aires, Argentina pedromemolibuffa@uba.ar          Luciano Del Corro ELIAS Lab, Departamento de Ingeniería Universidad de San Andrés Victoria, Argentina delcorrol@udesa.edu.ar
@@ 1 Introduction
Crucially, monitoring requires a quantity in accuracy units—not only a relative uncertainty score. Entropy- and logprob-based traces carry meaningful uncertainty signals (Malinin and Gales, 2021; Kuhn et al., 2023; Bouchard et al., 2025a; Kadavath et al., 2022), but raw scores are typically not directly interpretable as accuracy and can vary in scale across models and domains. We therefore cast monitoring as predicting correctness probabilities from decoding traces and aggregating them into domain-level accuracy estimates, yielding outputs that are directly actionable.
@@
We evaluate in a controlled STEM reasoning setting with ten benchmarks spanning elementary math, advanced math, and science. To stress-test robustness under domain shift, we exhaustively vary supervision: for each $k\in\{1,2,3,4\}$ we train on all $\binom{10}{k}$ benchmark subsets and estimate accuracy on the remaining 10k10-k. We repeat this across nine LLMs from six families (3B–20B) and multiple estimator variants (classifier family, calibration, class balancing), totaling >41,000>41{,}000 configurations. The resulting estimates track held-out benchmark accuracy closely and achieve high rank agreement (Figure 1); we report both ranking quality (Spearman $\rho$) and accuracy-level fit (AEE). Reliability varies across models and should be validated for the target serving model before deployment. [^1] Code for reproducing results can be found on this anonymous repository.
Overall, our results suggest that lightweight classifiers trained on entropy profiles provide a practical signal that can support both continuous monitoring and targeted data acquisition. We intentionally establish this in a STEM setting with verifiable correctness, enabling a thorough robustness study under a fixed protocol—exhaustive train/test benchmark compositions, multiple estimator variants, and multiple model families. Although our experiments use open-weight models, we restrict ourselves to top-$k$ decoding logprob signals (top-20) exposed by common model-serving APIs, keeping the approach compatible with both open and closed deployments. Extending to less verifiable, open-ended domains and validating transfer to closed commercial models are natural next steps.
@@ 2 From Signatures to Accuracy Estimates
Building on evidence that token-probability traces carry information about response correctness Malinin and Gales (2021); Kuhn et al. (2023); Ali et al. (2025); Bouchard et al. (2025a), we study whether a cheap signal available at inference time can support domain-level performance monitoring in STEM reasoning. We summarize the model’s output uncertainty signature as the sequence of token-level entropies computed from next-token probabilities during generation.
@@
Design requirements. For practical monitoring, the signal and estimator should be: (i) computationally lightweight (single-pass, logprob-only features), (ii) API-compatible (requiring only top-$k$ token probabilities, not hidden states), and (iii) robust to domain shift (trained on one labeled source benchmark and applied to unseen domains). Accordingly, we focus on final-layer output probabilities logged during standard decoding and evaluate generalization by training on a single source benchmark and estimating accuracy across diverse unseen STEM reasoning benchmarks.
@@ 3 Entropy Profile Signals
To satisfy requirements (i) efficiency and (ii) API compatibility, we restrict ourselves to uncertainty signals available from standard decoding logs. Concretely, we ask whether the output-entropy trajectory contains enough signal to discriminate correct from incorrect responses using only compact summaries. We approximate entropy from top-$k$ log-probabilities, setting k=20k{=}20 to match the maximum exposed by commercial APIs.
Do simple summaries carry correctness signal? For each response, we compute an entropy trajectory over decoding steps, $\{\tilde{H}^{(t)}\}_{t=1}^{T}$. Following standard evaluation practice in uncertainty quantification (Malinin and Gales, 2021; Kuhn et al., 2023; Bouchard et al., 2025a; Kadavath et al., 2022), we test whether single-number summaries of this trajectory already discriminate correct from incorrect outputs. Concretely, for each summary statistic $s(\{\tilde{H}^{(t)}\})$, we treat larger values as a higher score for incorrectness and report its AUROC. For context, we also include white-box uncertainty baselines computed from token log-probabilities (SEA, $\mathrm{NLL}_{\mathrm{avg}}$, $\mathrm{NLL}_{\mathrm{max}}$, $\mathrm{NLL}_{\mathrm{sum}}$, LNTP, MTP, and PPL), defined in Appendix A.
@@
Table 1, shows those uncertainity metrics for 3 representative datasets and models (more models in appendix B). The vast majority of summary statistics achieve AUROC >0.5>0.5 across model–benchmark pairs, indicating non-trivial separability between the entropy distributions of correct and incorrect generations (Figure 2). Values tend to be high for all the other logprobs metrics.
As already shown in previous work, this pattern reflects that incorrect responses tend to exhibit higher entropy (lower confidence), causing many of these statistics to behave as natural uncertainty scores where simple thresholding yields better-than-chance discrimination.
However, no single summary is reliably best: the top feature shifts across both models and benchmarks (Table 1). This variability suggests that different aspects of the entropy profile capture complementary uncertainty cues, and that a single scalar is unlikely to serve as a universal monitoring signal. For example, lower-tail quantiles (Q10/Q25) are most predictive on MATH, consistent with sustained uncertainty during long derivations, whereas dispersion (Std) is more informative on GSM8K. Simple extremes and accumulations are also strong: the entropy accumulation score (SEA) and $\text{NLL}_{\text{sum}}$ consistently rank among the best baselines, and Max entropy is competitive across several models (e.g., SEA=0.8762 for ministral-3 8B on MATH; Max=0.8137 for gpt-oss 20B), suggesting that total uncertainty mass and peak uncertainty events are particularly diagnostic of errors. Higher-order moments (skewness, kurtosis) help in some settings (e.g., gpt-oss on OlympiadBench; 0.8558/0.8607) but collapse toward chance in others (e.g., Gemma-3 12B on MATH). Overall, relying on any single metric risks missing domain- and model-specific failure modes, motivating our use of a compact multi-statistic profile.
Overall, these entropy- and logprob-derived metrics provide a real correctness signal: many separate correct from incorrect generations well above chance, and some even track domain difficulty via aggregated scores. However, taken in isolation they remain uncertainty scores, not accuracy estimates. Their scales are model- and domain-dependent, they are not calibrated to probabilities, and the best-performing statistic can change with the benchmark and model family. As a result, threshold choices and score magnitudes are hard to interpret operationally (e.g., an increase from 1.2 to 1.5 does not translate into “accuracy dropped by 8 points”). This motivates the next step: rather than selecting a single metric and treating it as performance, we learn a probabilistic mapping from a compact entropy-profile vector to instance-level correctness probabilities, which can then be averaged to produce slice-level accuracy estimates in the units required for monitoring.
@@
A compact discriminative signature. Motivated by these observations, we avoid committing to a single “best” scalar. Instead, we encode each response with a fixed 11D entropy-profile vector capturing central tendency/dispersion (max, mean, std), distributional tails (Q10–Q90), shape (skewness, kurtosis), and accumulation (SEA). A lightweight probabilistic classifier can then learn which aspects matter for correctness in a given setting (Sec. 4), without brittle manual feature selection. We restrict features to entropy-trajectory summaries to keep the representation compact, motivated by recent work Ali et al. (2025) that demostrated that entropy signatures by themselves can achieve remarkable performance regarding incorrect answer detection.
4 Accuracy Estimation via Entropy Profile Features
[…]
Probabilistic model. We map entropy features to an estimated probability of correctness with a lightweight probabilistic model $f:\mathbb{R}^{11}\rightarrow[0,1]$, defining $\hat{P}(x)=f(\mathbf{h}_{x})$, where $\hat{P}(x)$ estimates the probability a response is correct.
From instance probabilities to domain-level accuracy. Given per-instance correctness probabilities $\hat{P}(x)$, we estimate accuracy on a target domain (or slice) $D$ by averaging $\hat{P}(x)$ over instances in $D$ (Eq. 1). This estimator requires only decoding-time log-probabilities at test time.
5 Experimental Setup
@@
We test whether compact entropy-profile features support domain-level accuracy estimation under domain shift: train an instance-level correctness predictor on a small set of benchmarks, then estimate accuracy on unseen benchmarks by aggregating predicted correctness probabilities.
@@
Benchmarks. We evaluate on ten STEM reasoning benchmarks spanning math and science: GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), GSM-Symbolic Mirzadeh et al. (2025), MATH Hendrycks et al. (2021), TheoremQA Chen et al. (2023), SciBench Wang et al. (2023), MatSciBench Zhang et al. (2025b), OlympiadBench He et al. (2024), LiveMathBench Anonymous (2025), and GPQA Rein et al. (2023). All tasks use zero-shot chain-of-thought prompting with free-form final answers. For benchmarks originally multiple-choice (GPQA, SciBench), we remove answer options from the prompt. For OlympiadBench we restrict to text-only math and physics questions; for LiveMathBench we use the v202505_all_en subset. [^2] https://huggingface.co/datasets/opencompass/LiveMathBench When a split exists, we evaluate on the test portion; otherwise we evaluate on the full benchmark.
@@
Instance labeling and prediction target. Each benchmark instance provides a question $q$, a reference answer $y^{\star}$, and a model-generated response $\hat{y}$. We extract the model’s final answer from $\hat{y}$ using benchmark-specific post-processing (e.g., stripping formatting and selecting the last boxed/numeric expression when applicable). An external validator LLM (Grok-4.1-Fast-Reasoning xAI (2025)) receives $(q,\hat{y}_{\text{final}},y^{\star})$ and outputs a binary label $z\in\{0,1\}$ indicating whether the final answer matches the reference; a manual audit of 100 randomly sampled instances yielded $97\%$ agreement with human judgment. We treat $z$ as supervision for our probabilistic correctness predictor.
@@
Models. We evaluate nine LLMs (3B-20B–six families) , and restrict features to top-20 decoding logprobs, matching a common interface constraint in logprob-returning serving stacks. We run the full pipeline separately for Ministral-3 3B Mistral AI (2025), Phi-3.5-Mini 3.8B Microsoft (2024), Qwen-3 4B Qwen Team (2025), Gemma-3 4B Google DeepMind (2025), Qwen-3 8B, Ministral-3 8B, Llama-3.1 8B Meta AI (2024), Gemma-3 12B, and GPT-OSS 20B OpenAI (2025).
Estimators and ablations. We evaluate three classifiers on the 11D entropy-profile features: logistic regression with $\ell_{1}$ regularization, random forest, and a multilayer perceptron (MLP). For each classifier, we select hyperparameters via cross-validated grid search on the training group, optimizing ROC-AUC. We also vary two training choices: class balancing (on/off) and isotonic calibration (on/off). Across 9 models, 385 groups, 3 classifier families, and $\times2$ training options, this produces $9\times 385\times 3\times 4=41{,}580$ configurations.
Domain-level evaluation metrics. For each held-out benchmark/domain $D$, we estimate accuracy by averaging per-instance correctness probabilities (Eq. 1). We report: (i) accuracy estimation error (AEE), mean absolute error between estimated and true benchmark accuracies over held-out domains; and (ii) Spearman correlation $\rho$ between estimated and true accuracies, capturing whether the estimator ranks domains for data acquisition.
Implementation. All LLMs are served with vLLM Kwon et al. (2023) (seed 42) with a maximum generation length of 2,048. Temperature 0.5 has been recommended for Shannon Entropy in (Kuhn et al., 2023). Classifiers are implemented in scikit-learn Pedregosa et al. (2011); features are z-scored using training-group statistics. Further details for reproducibility are presented in the appendix C.

### v3

@@ Abstract
Deploying LLMs raises two coupled challenges: (1) monitoring—estimating where a model underperforms as traffic and domains drift—and (2) improvement—prioritizing data acquisition to close the largest performance gaps. We test whether an inference-time signal can estimate slice-level accuracy under domain shift. For each response, we compute an output-entropy profile from final-layer next-token probabilities (from top-$k$ logprobs) and summarize it with different statistics. A lightweight classifier predicts instance correctness, and averaging predicted probabilities yields a domain-level accuracy estimate. We evaluate on ten STEM reasoning benchmarks with exhaustive train/test compositions ($k\in\{1,2,3,4\}$; all $\binom{10}{k}$ combinations), on different classifier models and features across nine LLMs from six families (3B–20B). Estimates often track held-out benchmark accuracy, and several models show near-monotonic ordering of domains, providing evidence for output-entropy profiles being an accessible signal for scalable monitoring and for targeted data acquisition.
@@
Pedro Memoli Buffa Departamento de Matematica, FCEyN Universidad de Buenos Aires Buenos Aires, Argentina pedromemolibuffa@uba.ar          Luciano Del Corro ELIAS Lab, Departamento de Ingeniería Universidad de San Andres Victoria, Argentina delcorrol@udesa.edu.ar
@@ 1 Introduction
Crucially, monitoring requires a quantity in accuracy units—not only a relative uncertainty score. Entropy- and logprob-based traces carry meaningful uncertainty signals (Malinin and Gales, 2021; Kuhn et al., 2023; Bouchard and Chauhan, 2025a; Kadavath et al., 2022), but raw scores are typically not directly interpretable as accuracy and can vary in scale across models and domains. We therefore cast monitoring as predicting correctness probabilities from decoding traces and aggregating them into domain-level accuracy estimates, yielding outputs that are directly actionable.
@@
We evaluate in a controlled STEM reasoning setting with ten benchmarks spanning elementary math, advanced math, and science. To stress-test robustness under domain shift, we exhaustively vary supervision: for each $k\in\{1,2,3,4\}$ we train on all $\binom{10}{k}$ benchmark subsets and estimate accuracy on the remaining 10k10-k. We repeat this across nine LLMs from six families (3B–20B) and multiple estimator variants (classifier family, calibration, class balancing and feature set), totaling >160,000>160{,}000 configurations. The resulting estimates often track held-out benchmark accuracy closely and achieve high rank agreement, with some LLMs presenting almost perfect calibration on sensible configurations (Figure 1). We nevertheless found that reliability varies across models and should be validated for the target serving model before deployment. [^1] Code for the paper can be found here.
Overall, our results suggest that lightweight classifiers trained on entropy profiles provide a practical signal that can support both continuous monitoring and targeted data acquisition. We intentionally establish this in a STEM setting with verifiable correctness, enabling a thorough robustness study under a fixed protocol—exhaustive train/test benchmark compositions, multiple estimator and feature variants, and multiple model families. Although our experiments use open-weight models, we restrict ourselves to top-$k$ decoding logprob signals (top-20) exposed by common model-serving APIs, keeping the approach compatible with both open and closed deployments. Extending to less verifiable, open-ended domains and validating transfer to closed commercial models are natural next steps.
@@ 2 From Signatures to Accuracy Estimates
Previous work has shown that token-probability traces carry information about response correctness Malinin and Gales (2021); Kuhn et al. (2023); Ali et al. (2025); Bouchard and Chauhan (2025a). We build on this finding and propose a simple two-stage methodology for domain-level accuracy monitoring: (1) extract a compact uncertainty feature vector from the decoding trace of each response, and (2) train a lightweight probabilistic classifier to predict instance-level correctness, whose predictions are then aggregated into slice-level accuracy estimates.
@@
Design requirements. For practical monitoring, the signal and estimator should be: (i) computationally lightweight (single-pass, no auxiliary models), (ii) API-compatible (requiring only top-$k$ token probabilities, not hidden states or full-vocabulary distributions), and (iii) robust to domain shift (transferring from a small labeled source to unseen domains). In Section 3 we identify a compact feature space derived from the entropy trajectory that satisfies (i) and (ii), and empirically validate that it carries instance-level correctness signal. In Section 4 we describe the evaluation protocol designed to stress-test (iii) under different LLMs and configurations, and in Section 5 we present the results.
@@ 3 Entropy Profile Signals
To satisfy requirements (i) efficiency and (ii) API compatibility, we restrict ourselves to uncertainty signals available from standard decoding logs and evaluate whether the output-entropy trajectory contains enough signal to discriminate correct from incorrect responses using only compact summaries. We approximate entropy from top-$k$ log-probabilities, setting k=20k{=}20 to match the maximum exposed by commercial APIs.
Do simple summaries carry correctness signal? For each response, we compute an entropy trajectory over decoding steps, $\{\tilde{H}^{(t)}\}_{t=1}^{T}$. Following standard evaluation practice in uncertainty quantification (Malinin and Gales, 2021; Kuhn et al., 2023; Bouchard and Chauhan, 2025a; Kadavath et al., 2022), we test whether single-number summaries of this trajectory already discriminate correct from incorrect outputs. Concretely, for each summary statistic $s(\{\tilde{H}^{(t)}\})$, we treat larger values as a higher score for incorrectness and report its AUROC. We also include traditional white-box uncertainty metrics computed from token log-probabilities (SEA, $\mathrm{NLL}_{\mathrm{avg}}$, $\mathrm{NLL}_{\mathrm{max}}$, $\mathrm{NLL}_{\mathrm{sum}}$, LNTP, MTP, and PPL), defined in Appendix A.
@@
Table 1 reports AUROC for three representative model–benchmark pairs (additional models in Appendix B). Since each benchmark provides reference answers, we obtain ground-truth correctness labels for every generation and use them to evaluate how well each summary statistic separates correct from incorrect responses. The vast majority of statistics achieve AUROC >0.5>0.5, confirming non-trivial separability: correct responses tend to concentrate in lower-entropy regions, reflecting higher model confidence during decoding, while incorrect responses shift toward higher entropy values, indicating greater token-level uncertainty (Figure 2). This is consistent with previous findings and causes many of these statistics to behave as natural uncertainty scores, where simple thresholding yields better-than-chance discrimination.
@@
However, no single summary is reliably best: the top feature shifts across both models and benchmarks (Table 1), suggesting that different aspects of the entropy profile capture complementary uncertainty cues. For example, lower-tail quantiles (Q10/Q25) are most predictive on MATH, whereas dispersion (Std) is more informative on GSM8K. Accumulation metrics (SEA, $\text{NLL}_{\text{sum}}$) consistently rank among the strongest individual signals, while higher-order moments (skewness, kurtosis) help in some settings but collapse toward chance in others. This variability suggests that relying on any single metric risks missing domain- and model-specific failure modes, motivating our use of a compact multi-statistic profile.
@@
A compact discriminative signature. Motivated by these observations, we avoid committing to a single “best” scalar. Instead, we encode each response with a fixed 17D entropy-profile vector capturing central tendency and dispersion (max, mean, std), distributional tails (Q10–Q90), shape (skewness, kurtosis), accumulation (SEA), and the traditional white-box UQ metrics presented in Appendix A. A lightweight probabilistic classifier can then learn which aspects matter for correctness in a given setting, without brittle manual feature selection.
4 Robustness Evaluation Protocol
@@
We now evaluate whether the methodology supports domain-level accuracy estimation under domain shift: we train instance-level correctness predictors on a small set of labeled benchmarks and estimate accuracy on unseen benchmarks by aggregating predicted correctness probabilities.
@@
Benchmarks. We evaluate on eleven STEM reasoning benchmarks spanning math and science: GSM8K Cobbe and others (2021), SVAMP Patel and others (2021), GSM-Symbolic Mirzadeh and others (2025), MATH Hendrycks and others (2021), TheoremQA Chen and others (2023), SciBench Wang and others (2023), MatSciBench Zhang and others (2025), OlympiadBench He and others (2024), LiveMathBench Anonymous (2025), and GPQA Rein and others (2023). All tasks use zero-shot chain-of-thought prompting with free-form final answers. For benchmarks originally multiple-choice (GPQA, SciBench), we remove answer options from the prompt. For OlympiadBench we restrict to text-only math and physics questions; for LiveMathBench we use the v202505_all_en subset. [^2] https://huggingface.co/datasets/opencompass/LiveMathBench When a split exists, we evaluate on the test portion; otherwise we evaluate on the full benchmark.
@@
Instance labeling and prediction target. Each benchmark instance provides a question $q$, a reference answer $y^{\star}$, and a model-generated response $\hat{y}$. We extract the model’s final answer from $\hat{y}$ using benchmark-specific post-processing (e.g., stripping formatting and selecting the last boxed/numeric expression when applicable). An external validator LLM (Grok-4.1-Fast-Reasoning xAI (2025)) receives $(q,\hat{y}_{\text{final}},y^{\star})$ and outputs a binary label $z\in\{0,1\}$ indicating whether the final answer matches the reference; a manual audit of 300 randomly sampled instances yielded $97\%$ agreement with human judgment. We treat $z$ as supervision for our probabilistic correctness predictor.
Models. We evaluate nine LLMs (3B–20B, six families), restricting features to top-20 decoding logprobs to match a common interface constraint in logprob-returning serving stacks. We run the full pipeline separately for Ministral-3 3B Mistral AI (2025), Phi-3.5-Mini 3.8B Microsoft (2024), Qwen-3 4B Qwen Team (2025), Gemma-3 4B Google DeepMind (2025), Qwen-3 8B, Ministral-3 8B, Llama-3.1 8B Meta AI (2024), Gemma-3 12B, and GPT-OSS 20B OpenAI (2025).
Baselines The individual UQ metrics from Appendix A return unbounded uncertainty scores that cannot directly serve as accuracy estimates. To create a comparable baseline, we calibrate each metric via Platt scaling (training a logistic regression) over the training set, mapping its scores to correctness probabilities that can be aggregated via Eq. 1. This yields a single-feature, logistic regression variant of our pipeline for each metric.
Estimator configurations. We evaluate three classifier families: logistic regression, random forest, and a multilayer perceptron (MLP). For each classifier, we select hyperparameters via cross-validated grid search on the training group, optimizing ROC-AUC. We also vary two training choices: class balancing (on/off) and isotonic calibration (on/off).
To test whether the full 17D vector is necessary or whether simpler summaries suffice, we train each configuration over four feature subsets of decreasing size: (i) the full 17D profile, (ii) the 10 entropy-distribution statistics only, (iii) max, $\text{SE}_{\text{sum}}$, and $\text{NLL}_{\text{sum}}$—the three most consistently discriminative signals from the AUROC analysis in Section 3—and (iv) $\text{SE}_{\text{sum}}$ alone, the single strongest individual metric.
Across 9 models, 385 train/test groups, 3 classifier families, $\times2$ training options, 4 feature subsets, and calibrated single-metric baselines, this produces over 160,000 configurations.
Domain-level evaluation metrics. For each held-out benchmark $D$, we estimate accuracy by averaging per-instance correctness probabilities (Eq. 1). We report: (i) accuracy estimation error (AEE), the mean absolute error between estimated and true benchmark accuracies over held-out domains; and (ii) Spearman correlation $\rho$ between estimated and true accuracies, capturing whether the estimator correctly ranks domains for data-acquisition prioritization.
Implementation. All LLMs are served with vLLM Kwon et al. (2023) (seed 42) with a maximum generation length of 2,048 tokens and temperature 0.5, as recommended for Shannon entropy estimation in Kuhn et al. (2023). Classifiers are implemented in scikit-learn Pedregosa et al. (2011); features are z-scored using training-group statistics. Further reproducibility details are in Appendix C.
@@
Method Spearman $\rho$ $\uparrow$ AEE $\downarrow$
Default (RQ1 Config) .95.05 .08.04
$\text{NLL}_{\text{sum}}$ .96.07 .10.04
$\text{SE}_{\text{sum}}$ .94.08 .10.05
$\text{SE}_{\text{max}}$ .95.07 .08.07
$\text{NLL}_{\text{max}}$ .90.43 .22.15
$\text{SE}_{\text{mean}}$ .84.21 .17.10
MTP .90.42 .23.14
$\text{NLL}_{\text{avg}}$ .77.23 .21.07
LNTP .61.50 .23.14
PPL .77.23 .21.06
Table 3: Median (IQR as subscript) across 9 LLMs under the Extremes training group.
5 Results
Research Questions. We organize results around four questions: (RQ1) does the entropy-profile signal support cross-domain accuracy estimation under plausible defaults? (RQ2) how does it compare with simple baselines? (RQ3) how sensitive is performance to which benchmarks are used for supervision? (RQ4) how sensitive is performance to estimator design choices (classifier family, calibration, balancing, feature subset)?
RQ1: Accuracy estimation under deployment-plausible defaults.
We report results for two training groups chosen a priori to reflect plausible supervision strategies rather than tuned for best score: Extremes (GSM8K + OlympiadBench), which pairs an elementary and a competition-level benchmark to span the widest difficulty range, and Intermediate (MATH + SciBench), which covers mid-range difficulty. Both use a random forest with class balancing, isotonic calibration, and the 10 entropy-distribution statistics as the feature subset. Table 2 and Figure 3 report cross-domain accuracy estimation quality.
Both training groups generalize out of domain for most LLMs, but Extremes is consistently stronger: seven of nine models achieve $\rho\geq 0.90$ with low error (AEE 0.030.030.120.12), while Intermediate yields systematically higher error and weaker ordering across all models (AEE 0.060.060.170.17). The signal is also model-dependent: Phi-3.5-Mini achieves near-perfect ordering (AEE $0.03$, $\rho=1.00$), whereas Qwen3-8B exhibits weaker agreement in both settings (AEE 0.120.120.170.17, $\rho\approx 0.75$). Additional training groups and estimator variants are reported in Appendix D.
RQ2: Comparison with single-metric baselines.
Having established that the entropy-profile estimator generalizes well under plausible defaults, we now ask whether this requires combining multiple features into a richer classifier, or whether a single calibrated UQ metric suffices. We compare the RQ1 estimator against the single-metric Platt-scaled baselines as described in Section 4, trained on the same Extremes group. Table 3 summarizes median performance and variability across the nine LLMs.
Three baselines are remarkably performant: $\text{NLL}_{\text{sum}}$, $\text{SE}_{\text{sum}}$, and $\text{SE}_{\text{max}}$ achieve $\rho$ and AEE in the same ballpark as the default estimator, with the default being only slightly more consistent across LLMs. The remaining baselines degrade considerably—median $\rho$ drops to 0.610.610.900.90 with large variability (IQR up to $0.50$), and median AEE roughly doubles. This separation is consistent with the AUROC analysis in Section 3, where accumulation and extrema metrics were also the strongest individual discriminators of correctness. Overall, the narrow gap between the top baselines and the multi-feature estimator suggests that training composition and having accumulation or extrema metrics within the feature set is far more important than classifier architecture or feature dimensionality. We evaluate this exhaustively in the following sections.
RQ3: Sensitivity to training composition.
We quantify how accuracy estimation depends on the composition of the supervision set. Table 4 aggregates results over all training groups $G$ of size $k\in\{1,2,3,4\}$ and over classifier configurations, reporting median AEE and its IQR. Increasing $k$ produces two consistent effects across all nine LLMs: median error decreases monotonically, and sensitivity to benchmark choice shrinks sharply. At k=1k{=}1, median AEE ranges from $0.20$ to $0.27$ with high variability; by k=4k{=}4, it improves to 0.060.060.160.16 with IQR compressed to 0.030.030.040.04. Benchmark choice is thus a first-order design decision at small $k$, but becomes less critical as supervision spans more tasks.
Difficulty balance explains much of the composition effect. To make this dependence interpretable, we summarize each training group $G$ by its weighted average accuracy—the mean accuracy $A(D)$ over benchmarks $D\in G$ and LLMs, weighted by instance counts in the group—and relate it to held-out estimation quality. We perform this under the RQ1 configuration to control for architecture variability. Figure 4 reveals a U-shaped relationship: supervision sets that are too easy or too hard generalize worse, whereas groups with intermediate weighted accuracy (roughly 0.40.40.70.7) yield the lowest AEE. This is consistent with difficulty-diverse groups exposing the estimator to both low-entropy success patterns and high-entropy failure patterns, while homogeneous groups underrepresent one regime and miscalibrate under domain shift. We analyze the best and worst benchmark combinations and a leave-one-out sensitivity study in Appendix F.